-
CycleTrack and SDN are teaming up to make medical school admissions more transparent and accessible! Read the announcement to learn how we’re supporting the future of CycleTrack’s free application-tracking tools .
You are using an out of date browser. It may not display this or other websites correctly.
You should upgrade or use an alternative browser.
You should upgrade or use an alternative browser.
Get help with your application
Use all the free resources available to you from SDN: articles, guides, expert advising, forums discussions, and school research.
A container holds 8 red and 4 white balls. Two balls are drawn in sequence without replacement from the container. Which of the following represents the probability that exactly one of these balls is red?
if first ball is red, (8/12)*(4/11)
if second ball is read, (4/12)*(8/11)
add them up you got 6/11?
there is 1/2 chance for 1st ball being red, and same 1/2 chance for 2nd ball being red.
So instead of simply adding up (8/12)*(4/11) + (4/12)*(8/11), you should do (1/2)(8/12)*(4/11)+ (1/2)(4/12)*(8/11) = 32/132
So instead of simply adding up (8/12)*(4/11) + (4/12)*(8/11), you should do (1/2)(8/12)*(4/11)+ (1/2)(4/12)*(8/11) = 32/132
Advertisement - Members don't see this ad
The guy way up there got it right (just added them incorrectly). No need to keep chiming in with wrong answers. The correct answer is A.
Wrong. There is not a 1/2 chance you'll draw a red ball when there are clearly DOUBLE the amount of red balls to start.
Wrong answer. The question asked for 1 of each. You solved for 2 reds.
there is 1/2 chance for 1st ball being red, and same 1/2 chance for 2nd ball being red.
So instead of simply adding up (8/12)*(4/11) + (4/12)*(8/11), you should do (1/2)(8/12)*(4/11)+ (1/2)(4/12)*(8/11) = 32/132
Wrong. There is not a 1/2 chance you'll draw a red ball when there are clearly DOUBLE the amount of red balls to start.
excuse me if this is a wrong answer, but can anyone tell me if this is right/wrong?
add up 8 red and 4 white = 12 total
Draw 1 (red): 8/12
Draw 2 (red): 7/11
P(red) = (8/12)*(7/11) = 56/132 - answer D
Wrong answer. The question asked for 1 of each. You solved for 2 reds.
Streetwolf. yup you are right. my mistake. no half half.
you have to calculate for both balls then add together,
8/12 * 4/11 = 8/33 next 4/12 * 8/11 = 8/33 now add together you get 16/33 answer A.
This is from 2010.
😱This is from 2010.